# Supervised heterogeneous network estimation via survival-based Bayesian graphical models

### v2

@@ 2.5 Theoretical properties
Suppose that Assumptions 1-5 hold. Let $\widetilde{s}=d_{\bm{{\beta}}}\sqrt{d_{\bm{{\mu}}}}+d_{\bm{{\beta}}}^{2}+\sqrt{p}+\sqrt{{s+p/K}}$. Assume that $\widetilde{s}\sqrt{{K^{3}\log p}}=o(\sqrt{n})$ and the lower bound of the minimal signal in the true parameters satisfies $\min\limits_{k=1,\ldots,K}\left\{\min\limits_{(j,l)\in\mathcal{W}_{k}^{*}}|\omega_{k,jl}^{*}|,\min\limits_{j\in\mathcal{U}_{k}^{*}}|\mu_{k,j}^{*}|,\min\limits_{j\in\mathcal{V}_{k}^{*}}|\beta_{k,j}^{*}|\right\}>2\widetilde{s}\sqrt{{K^{3}\log p}/{n}}$. Then,
1. The non-asymptotic bound of the estimation error of $\hat{\bm{{\Psi}}}$ obtained with the proposed EM algorithm is
$\left\|\hat{\bm{\Psi}}-\bm{\Psi}^{*}\right\|_{2}=O_{p}\left(\widetilde{s}\sqrt{{K^{3}\log p}/{n}}\right).$ (3)
2. For the final estimators $\widetilde{\bm{\Omega}}_{k}$, $\widetilde{\bm{\mu}}_{k}$, and $\widetilde{\bm{\beta}}_{k}$ with $\widetilde{\omega}_{k,jl}=\hat{\omega}_{k,jl}{I}\left\{\left|\hat{\omega}_{k,jl}\right|>\widetilde{s}\sqrt{{K^{3}\log p}/{n}}\right\}$, $\widetilde{\mu}_{k,j}=\hat{\mu}_{k,j}{I}\left\{\left|\hat{\mu}_{k,j}\right|>\widetilde{s}\sqrt{{K^{3}\log p}/{n}}\right\}$, and $\widetilde{\beta}_{k,j}=\hat{\beta}_{k,j}{I}\left\{\left|\hat{\beta}_{k,j}\right|>\widetilde{s}\sqrt{{K^{3}\log p}/{n}}\right\}$, we have $\left\|\widetilde{\bm{\Omega}}-\bm{\Omega}^{*}\right\|_{2}=O_{p}\left(\widetilde{s}\sqrt{{K^{3}\log p}/{n}}\right)$ and with probability approaching 1, $\widehat{\mathcal{W}}_{k}=\mathcal{W}_{k}^{*}$, $\widehat{\mathcal{U}}_{k}=\mathcal{U}_{k}^{*}$, and $\widehat{\mathcal{V}}_{k}=\mathcal{V}_{k}^{*}$ for any $k=$ $1,\ldots,K$, where $\widetilde{\bm{\Omega}}=\operatorname{vec}\left(\widetilde{\bm{\Omega}}_{1},\ldots,\widetilde{\bm{\Omega}}_{K}\right)$, and $\widehat{\mathcal{W}}_{k}$ denotes the set of the nonzero off-diagonal elements of $\widetilde{\bm{\Omega}}_{k}$, and $\widehat{\mathcal{U}}_{k}$ and $\widehat{\mathcal{V}}_{k}$ denote the sets of the nonzero elements of $\widetilde{\bm{\mu}}_{k}$ and $\widetilde{\bm{\beta}}_{k}$, respectively.
In Theorem 1, if $K$ is fixed, the non-asymptotic bound of the estimation error reduces to $\left\|\widetilde{\bm{\Omega}}-\bm{\Omega}^{*}\right\|_{2}=O_{p}\left(\left(d_{\bm{{\beta}}}\sqrt{d_{\bm{{\mu}}}}+d_{\bm{{\beta}}}^{2}+\sqrt{{s+p}}\right)\sqrt{\log p/n}\right)$. Denote by $\mathcal{\bm{Q}}\equiv\left\{\bm{\Omega}=\operatorname{vec}\left(\bm{\Omega}_{1},\ldots,\right.\right.$ $\left.\left.\bm{\Omega}_{K}\right):\right.$ $\left.\bm{\Omega}_{k}\text{ is positive definite and~{}}\max_{k}\|\bm{\Omega}_{k}\|_{\infty}\leq C_{\mathcal{Q}}\right\}$ the class of precision matrices. Under Assumption 3, the upper bound $C_{\mathcal{Q}}$ of $\mathcal{\bm{Q}}$ is independent of $(n,p)$. Provided that $d_{\bm{{\beta}}}\sqrt{d_{\bm{{\mu}}}}+d_{\bm{{\beta}}}^{2}=O(\sqrt{s+p})$, the convergence rate of the precision matrix estimate $\left\|\widetilde{\bm{\Omega}}-\bm{\Omega}^{*}\right\|_{2}=O_{p}(\sqrt{(s+p)\log p/n})$, which achieves the minimax optimal rate for estimating an $s$-sparse precision matrix under the Frobenius norm as shown in [3]. The proof of Theorem 1 is provided in Section S5 of the supplementary materials.
@@ 3 Simulation
Simulation studies are conducted to evaluate the performance of the proposed approach under the following settings. (a) $K=2$ and $3$ and $p=100$. (b) Consider two settings for the sample size. One is a balanced design where there are 150 subjects in each subgroup, while the other is an imbalanced design with sample sizes being $100$ and $200$ for $K=2$ and $100,150$, and $200$ for $K=3$. Thus, there are $n=300$ subjects for $K=2$, and $n=450$ for $K=3$ in total. (c) The predictor observations $\left\{\bm{x}_{i}\right\}$ in the $k$th subgroup are simulated from $\mathcal{N}\left(\bm{0},\bm{\Omega}_{k}^{-1}\right)$, where $\bm{\Omega}_{k}$ is generated based on the corresponding network structure. Specifically, following [7], we simulate each network with ten unconnected subnetworks. To comprehensively examine the performance of the proposed approach under different degrees of similarity across subgroup-specific precision matrices, we consider three specific settings S1-S3, with the number of subnetworks sharing the same sparsity structure in all subgroups being 3, 5, and 7, respectively. For each subnetwork, we consider three types of structure, including the power-law, nearest-neighbor, and Erdös-Rényi networks. The detailed settings are provided in Section S6 of the supplementary materials. (d) Each logarithmic survival observation $t_{i}$ is generated as follows. First, the logarithmic survival times of the $k$th subgroup are generated from $\mathcal{N}\left(\bm{\beta}_{k}^{\top}\bm{x}_{i},0.01^{2}\right)$, where five elements of $\bm{\beta}_{k}$ are non-zero, and the rest $p-5$ elements are all zero with $\bm{\beta}_{1}=\left(2,2,2,2,2,0,\ldots,0\right)^{\top}$, $\bm{\beta}_{2}=-\left(2,2,2,2,2,0,\ldots,0\right)^{\top}$, and $\bm{\beta}_{3}=\left(0,0,1,1,1,1,1,0,\ldots,0\right)^{\top}$, respectively. Then, the censoring time is simulated from a Gamma distribution with the shape and scale parameters controlling the censoring rate to be $\sim 20\%$. There are a total of 36 scenarios.
In addition to the proposed approach, six alternative approaches are also considered. BCGS is a Bayesian approach for simultaneously clustering and learning multiple sparse graphical models using a joint spike-and-slab graphical lasso prior [6]. SCAN is a mixture Gaussian graphical model-based approach that applies the group lasso penalty to accommodate common structures across unknown heterogeneous subgroups [12]. TGemBag and TJGL are two “ideal approaches” which assume that the true subgroup memberships are known and adopt GemBag and JGL for jointly estimating multiple networks for the true subgroups, respectively. Here, GemBag is a Bayesian graphical model for accommodating similarity as well as difference of multiple networks [28], and JGL employs a likelihood-based joint graphical Lasso for multiple network estimations [7]. FMGemBag and FMJGL are two two-stage approaches, where in the first stage, a finite mixture (FM) of AFT models developed in [24] is used to obtain the subgroup memberships of all subjects, which can account for right censoring and unknown heterogeneous predictor-survival associations, and in the second stage, GemBag and JGL are adopted for network estimation based on the estimated subgroup memberships, respectively. Among these methods, BCGS, SCAN, TJGL, and FMJGL can be realized using or extending the functionalities provided by the R package JGL, and TGemBag and FMGemBag can be realized using R codes at https://github.com/xinming104/GemBag. BCGS and SCAN achieve identification of unknown heterogeneity and network structures simultaneously, however, are unsupervised and only account for predictors. GemBag and JGL are developed for estimating precision matrices with known subgroup memberships. We consider two types of subgroup memberships for GemBag and JGL, including the one based on the FM approach, which utilizes the information of survival for heterogeneous analysis, and the “ideal” type based on the true subgroup memberships.
@@
To evaluate the performance of different approaches, we consider the clustering error (CE) for evaluating the heterogeneity identification performance, precision matrix squared error (PME) for evaluating the estimation performance, and true and false positive rates (TPR and FPR) for evaluating the network identification performance. The details are provided in Section S6 of the supplementary materials. Here, similar to other network-based heterogeneity analysis studies [12], we focus on the performance of estimation and identification on $\omega_{k,jl}$’s, which can be more challenging than on $\mu_{k,j}$’s and $\beta_{k,j}$’s.
@@
100 replicates are conducted for each scenario, where the true value of $K$ is used for all approaches. The summary results for $K=2$ under power-law and Erdös-Rényi networks are presented in Tables 2-3, and the rest of the results are provided in Section S6 of the supplementary materials. Under all simulated scenarios, the proposed approach consistently demonstrates superior or competitive performance compared to the alternative approaches. In terms of subgroup identification, the proposed approach performs significantly better with much smaller CEs. Among the alternatives, the FM methods tend to exhibit the poorest subgroup identification performance, while SCAN and BCGS have better subgroup identification accuracy but are still much inferior compared to the proposed approach.
@@
Table 2: Simulation results for the power-law network under the scenarios with $K=2$. In each cell, mean (SD) based on 100 replicates.
proposed BCGS SCAN TGemBag TJGL FMGemBag FMJGL
S1 with the balanced design
CE 0.001(0.003) 0.044(0.134) 0.045(0.134) - - 0.196(0.116) 0.196(0.116)
PME 4.502(0.362) 5.060(1.793) 5.054(1.784) 3.870(0.507) 7.789(0.528) 7.690(2.508) 8.695(1.577)
TPR 0.932(0.039) 0.901(0.142) 0.897(0.142) 0.933(0.043) 0.931(0.048) 0.871(0.128) 0.887(0.098)
FPR 0.025(0.005) 0.032(0.008) 0.031(0.008) 0.024(0.006) 0.025(0.008) 0.122(0.018) 0.080(0.022)
S2 with the balanced design
CE 0.002(0.005) 0.087(0.169) 0.088(0.170) - - 0.263(0.119) 0.263(0.119)
PME 4.398(0.335) 6.847(10.771) 5.453(2.296) 3.850(0.539) 7.821(0.583) 7.609(1.913) 8.581(1.299)
TPR 0.936(0.039) 0.870(0.174) 0.863(0.178) 0.932(0.051) 0.924(0.056) 0.866(0.116) 0.889(0.079)
FPR 0.025(0.006) 0.031(0.009) 0.030(0.008) 0.023(0.006) 0.024(0.009) 0.117(0.028) 0.081(0.032)
S3 with the balanced design
CE 0.009(0.010) 0.190(0.226) 0.195(0.230) - - 0.293(0.132) 0.293(0.132)
PME 4.134(0.348) 7.335(4.777) 6.917(3.424) 4.479(1.614) 8.222(1.080) 6.343(1.715) 8.367(3.251)
TPR 0.951(0.027) 0.775(0.255) 0.779(0.232) 0.867(0.172) 0.894(0.097) 0.910(0.123) 0.931(0.070)
FPR 0.023(0.006) 0.033(0.010) 0.033(0.009) 0.021(0.009) 0.022(0.011) 0.114(0.019) 0.074(0.022)
S1 with the imbalanced design
CE 0.001(0.002) 0.015(0.042) 0.015(0.039) - - 0.248(0.047) 0.248(0.047)
PME 4.929(0.446) 5.131(0.684) 5.147(0.685) 4.299(0.630) 7.806(0.514) 8.862(2.024) 9.379(1.350)
TPR 0.897(0.047) 0.877(0.084) 0.870(0.085) 0.887(0.065) 0.926(0.052) 0.803(0.122) 0.841(0.102)
FPR 0.027(0.006) 0.031(0.008) 0.031(0.008) 0.024(0.006) 0.029(0.009) 0.132(0.020) 0.091(0.021)
S2 with the imbalanced design
CE 0.005(0.009) 0.077(0.140) 0.078(0.140) - - 0.208(0.066) 0.208(0.066)
PME 4.713(0.381) 6.938(5.715) 7.252(6.652) 4.282(0.570) 7.838(0.587) 7.081(1.942) 8.155(1.178)
TPR 0.912(0.045) 0.830(0.165) 0.820(0.163) 0.889(0.061) 0.922(0.056) 0.884(0.104) 0.911(0.067)
FPR 0.025(0.007) 0.031(0.009) 0.030(0.009) 0.023(0.007) 0.028(0.010) 0.120(0.021) 0.086(0.027)
S3 with the imbalanced design
CE 0.013(0.009) 0.158(0.174) 0.162(0.176) - - 0.274(0.120) 0.274(0.120)
PME 4.449(0.381) 7.160(4.339) 7.243(4.615) 4.958(1.616) 8.249(1.091) 6.989(1.190) 7.981(0.816)
TPR 0.932(0.030) 0.775(0.192) 0.765(0.194) 0.817(0.171) 0.886(0.099) 0.877(0.089) 0.920(0.044)
FPR 0.022(0.006) 0.033(0.010) 0.032(0.009) 0.021(0.009) 0.025(0.012) 0.115(0.020) 0.083(0.017)
@@
Table 3: Simulation results for the Erdös-Rényi network under the scenarios with $K=2$. In each cell, mean (SD) based on 100 replicates.
proposed BCGS SCAN TGemBag TJGL FMGemBag FMJGL
S1 with the balanced design
CE 0.005(0.012) 0.027(0.072) 0.027(0.073) - - 0.193(0.071) 0.193(0.071)
PME 3.513(0.245) 3.658(0.861) 3.634(0.857) 2.919(0.388) 5.146(0.685) 7.077(2.002) 6.940(1.593)
TPR 0.962(0.036) 0.938(0.083) 0.935(0.083) 0.954(0.046) 0.947(0.048) 0.769(0.164) 0.766(0.149)
FPR 0.024(0.005) 0.027(0.007) 0.027(0.006) 0.024(0.006) 0.020(0.006) 0.140(0.031) 0.069(0.020)
S2 with the balanced design
CE 0.008(0.012) 0.059(0.108) 0.056(0.100) - - 0.218(0.116) 0.218(0.116)
PME 3.477(0.240) 3.807(1.067) 3.764(1.001) 2.851(0.380) 5.130(0.629) 5.589(1.822) 5.996(1.695)
TPR 0.971(0.032) 0.935(0.103) 0.932(0.097) 0.963(0.043) 0.952(0.046) 0.879(0.149) 0.873(0.136)
FPR 0.028(0.008) 0.030(0.008) 0.029(0.008) 0.026(0.008) 0.023(0.008) 0.132(0.037) 0.067(0.024)
S3 with the balanced design
CE 0.023(0.030) 0.135(0.172) 0.142(0.176) - - 0.251(0.142) 0.251(0.142)
PME 3.199(0.239) 4.276(1.707) 4.373(1.973) 2.768(0.296) 4.956(0.627) 5.157(1.151) 5.497(1.145)
TPR 0.975(0.023) 0.895(0.154) 0.886(0.156) 0.970(0.032) 0.960(0.038) 0.912(0.079) 0.916(0.081)
FPR 0.024(0.008) 0.030(0.009) 0.029(0.008) 0.026(0.008) 0.022(0.009) 0.139(0.033) 0.074(0.031)
S1 with the imbalanced design
CE 0.005(0.008) 0.033(0.070) 0.034(0.071) - - 0.230(0.092) 0.230(0.092)
PME 3.787(0.379) 4.062(0.932) 4.055(0.951) 3.253(0.582) 5.185(0.681) 6.182(1.978) 6.591(2.380)
TPR 0.933(0.051) 0.884(0.108) 0.880(0.107) 0.904(0.072) 0.938(0.050) 0.813(0.199) 0.819(0.191)
FPR 0.027(0.007) 0.028(0.007) 0.027(0.007) 0.025(0.007) 0.024(0.008) 0.130(0.032) 0.074(0.022)
S2 with the imbalanced design
CE 0.009(0.014) 0.064(0.114) 0.057(0.099) - - 0.195(0.098) 0.195(0.098)
PME 3.689(0.337) 4.200(1.240) 4.140(1.131) 3.122(0.517) 5.186(0.655) 6.258(1.715) 6.543(1.687)
TPR 0.951(0.042) 0.893(0.111) 0.888(0.107) 0.925(0.070) 0.947(0.056) 0.829(0.142) 0.838(0.124)
FPR 0.026(0.008) 0.030(0.009) 0.029(0.009) 0.026(0.008) 0.027(0.010) 0.123(0.025) 0.071(0.016)
S3 with the imbalanced design
CE 0.028(0.030) 0.142(0.149) 0.145(0.150) - - 0.265(0.116) 0.265(0.116)
PME 3.416(0.241) 5.082(2.511) 5.170(2.652) 3.072(0.373) 5.007(0.624) 4.896(0.902) 5.317(0.813)
TPR 0.962(0.026) 0.825(0.163) 0.812(0.162) 0.930(0.058) 0.958(0.036) 0.940(0.054) 0.947(0.060)
FPR 0.023(0.007) 0.031(0.009) 0.030(0.009) 0.027(0.008) 0.026(0.010) 0.131(0.033) 0.076(0.028)
As the similarity between the networks increases (from S1 to S3), it is more challenging to identify subgroup memberships. However, the proposed approach still shows more prominent advantages compared to SCAN, BCGS, FMGemBag, and FMJGL in terms of both heterogeneity and network identification performance. It is interesting that, under S3, although with a certain level of clustering errors, the proposed approach has better network identification performance than TGemBag and TJGL with “ideal” true subgroup memberships, suggesting the effectiveness of the proposed similarity-based prior. Under the scenarios with an imbalanced design or a larger number of subgroups, performance of all approaches declines. Under these scenarios, the proposed approach, when compared to TGemBag, exhibits a slightly worse performance in network identification and estimation due to the more complex unknown heterogeneity and the relatively limited sample size. However, it still consistently outperforms the other alternatives. It maintains its superiority under the scenarios with various types of network structure, including the more challenging power-law networks with denser structures and hub structures. With relatively simpler nearest neighbor networks, when all methods generally perform well with higher TPR values, the proposed approach retains significantly fewer FPs compared to most alternatives.
@@
Additionally, to gain a deeper understanding of the similarity-based prior, we illustrate a representative replicate under the scenario with setting S3 and an imbalanced design. The heatmaps depicting the true sparsity structures and the estimated results of the precision matrices with our proposed approach, SCAN, and TGemBag, are presented in Figures S3-S4 (supplementary materials). The findings indicate that our proposed approach excels in accurately identifying true sparsity structures and more effectively accommodating the similarity structures across different networks.
@@
Following previous studies [11, 12, 31], the main simulations assume that the number of subgroups $K$ is known for a more fair comparison. When $K$ is unknown, we recommend using the BIC for model selection. As an illustration, we consider scenarios with a power-law network and balanced design and evaluate the empirical performance of BIC over 100 replicates with candidate values $K=1,\ldots,5$. The selection frequencies, summarized in Table S5 (supplementary materials), indicate that BIC generally performs well. In particular, the S3 configuration, which exhibits greater subgroup similarity than S1 and S2, poses more challenges for accurate identification. For the BIC-selected values of $K$, Table S6 (supplementary material) reports the corresponding identification and estimation results. In cases where $K$ is incorrectly estimated, label alignment is performed minimizing the estimation error of $\bm{\Psi}$ [22]. The results remain consistent with those presented in Table 2 and Table S2 (supplementary materials).

### v4

@@ 3 Theoretical properties
Let $\tilde{s}=d_{\boldsymbol{{\beta}}}\sqrt{d_{\boldsymbol{{\mu}}}}+d_{\boldsymbol{{\beta}}}^{2}+\sqrt{p}+\sqrt{{s+p/K}}$. Assume that $\tilde{s}\sqrt{{K^{3}\log p}}=o(\sqrt{n})$. Then,
1. Under Assumptions 1-5 (supplementary materials), the non-asymptotic bound of the estimation error of $\hat{\boldsymbol{{\Psi}}}$ obtained with the proposed EM algorithm is
$\left\|\hat{\boldsymbol{\Psi}}-\boldsymbol{\Psi}^{*}\right\|_{2}=O_{p}\left(\tilde{s}\sqrt{{K^{3}\log p}/{n}}\right).$ (3)
2. Under Assumptions 1-6 (supplementary materials), the thresholded estimators $\tilde{\boldsymbol{\Theta}}=\hat{\boldsymbol{\Theta}}\cdot I\left\{|\hat{\boldsymbol{\Theta}}|>\tilde{s}\sqrt{K^{3}\log p/n}\right\}$ satisfies that $\left\|\tilde{\boldsymbol{\Omega}}-\boldsymbol{\Omega}^{*}\right\|_{2}=O_{p}\left(\tilde{s}\sqrt{{K^{3}\log p}/{n}}\right)$ and
$\mathbb{P}\left(\tilde{\gamma}_{k,jl}={\gamma}_{k,jl}\right)\rightarrow 1,\quad\mathbb{P}\left(\widehat{\mathcal{W}}_{k}={\mathcal{W}}_{k},\widehat{\mathcal{U}}_{k}=\mathcal{U}_{k}^{\star},\widehat{\mathcal{V}}_{k}=\mathcal{V}_{k}^{\star}\right)\rightarrow 1$
for any $k=$ $1,\ldots,K$ and $1\leq j<l\leq p$, where $\widehat{\mathcal{W}}_{k}=\{(j,l)\mid\tilde{\gamma}_{k,jl}>a\}$ with a threshold $a\in(0,1)$, and $\widehat{\mathcal{U}}_{k}$ and $\widehat{\mathcal{V}}_{k}$ denote the sets of the nonzero elements of $\tilde{\boldsymbol{\mu}}_{k}$ and $\tilde{\boldsymbol{\beta}}_{k}$, respectively.
[…]
We establish theoretical guarantees for estimation consistency in a challenging setting characterized by double-level heterogeneity, censored outcomes, and a novel similarity prior—a regime not covered by existing theory. Building on this foundation, we prove the consistency of the MAP plug-in PIPs, thereby providing a rigorous basis for network structure recovery and the quantification of Bayesian edge-connection uncertainty within our framework. This distinguishes our theoretical contributions from prior work: for instance, Gan et al. (2019) analyze a convex single-network model, while Hao et al. (2018) study an unsupervised multi-network mixture without outcome-driven heterogeneity. In contrast, our results simultaneously address non-convexity and supervised, outcome-dependent heterogeneity, offering novel theoretical support for network construction in this complex setting.
@@ 4 Simulation
Simulation studies are conducted to evaluate the performance of the proposed approach under the following settings. (a) $K=2$ and $3$ and $p=100$. (b) Consider two settings for the sample size. One is a balanced design where there are 150 subjects in each subgroup, while the other is an imbalanced design with sample sizes being $100$ and $200$ for $K=2$ and $100,150$, and $200$ for $K=3$. Thus, there are $n=300$ subjects for $K=2$, and $n=450$ for $K=3$ in total. (c) The predictor observations $\left\{\boldsymbol{x}_{i}\right\}$ in the $k$th subgroup are simulated from $\mathcal{N}\left(\boldsymbol{0},\boldsymbol{\Omega}_{k}^{-1}\right)$, where $\boldsymbol{\Omega}_{k}$ is generated based on the corresponding network structure. Specifically, following Danaher et al. (2014), we simulate each network with ten unconnected subnetworks. To comprehensively examine the performance of the proposed approach under different degrees of similarity across subgroup-specific precision matrices, we consider three specific settings S1-S3, with the number of subnetworks sharing the same sparsity structure in all subgroups being 3, 5, and 7, respectively. For each subnetwork, we consider three types of structure, including the power-law, nearest-neighbor, and Erdös-Rényi networks. The detailed settings are provided in Section S6 of the supplementary materials. (d) Each logarithmic survival observation $t_{i}$ is generated as follows. First, the logarithmic survival times of the $k$th subgroup are generated from $\mathcal{N}\left(\boldsymbol{\beta}_{k}^{\top}\boldsymbol{x}_{i},0.01^{2}\right)$, where five elements of $\boldsymbol{\beta}_{k}$ are non-zero, and the rest p5p-5 elements are all zero with $\boldsymbol{\beta}_{1}=\left(2,2,2,2,2,0,\ldots,0\right)^{\top}$, $\boldsymbol{\beta}_{2}=-\left(2,2,2,2,2,0,\ldots,0\right)^{\top}$, and $\boldsymbol{\beta}_{3}=\left(0,0,1,1,1,1,1,0,\ldots,0\right)^{\top}$, respectively. Then, the censoring time is simulated from a Gamma distribution with the shape and scale parameters controlling the censoring rate to be $\sim 20\%$. There are a total of 36 scenarios.
In addition to the proposed approach SBJGM, we compare six alternative approaches: BCGS (Dai and Jin, 2022), SCAN (Hao et al., 2018); two “ideal” methods, TGemBag and TJGL, which assume true subgroup memberships and apply group estimation of multiple Bayesian graphical models (GemBag, Yang et al. (2021)) and joint graphical lasso (JGL, Danaher et al. (2014)) to estimate subgroup-specific networks; and two two-stage methods, FMGemBag and FMJGL, which first use the finite-mixture AFT model (FM) of Shokoohi et al. (2019) to estimate subgroup memberships before applying GemBag and JGL. Details are provided in Section S6 of the supplementary materials.
@@
To evaluate the performance of different approaches, we consider the clustering error (CE) for evaluating the heterogeneity identification performance, precision matrix squared error (PME) for evaluating the estimation performance, and true and false positive rates (TPR and FPR) for evaluating the network identification performance. The details are provided in Section S6 of the supplementary materials. Here, similar to other network-based heterogeneity analysis studies (Hao et al., 2018), we focus on the performance of estimation and identification on $\omega_{k,jl}$’s, which can be more challenging than on $\mu_{k,j}$’s and $\beta_{k,j}$’s.
@@
100 replicates are conducted for each scenario, where the true value of $K$ is used for all approaches. The summary results for $K=2$ under power-law and Erdös-Rényi networks are presented in Figures 1 and 2, and the rest of the results are provided in Section S6 of the supplementary materials. Under all simulated scenarios, the proposed SBJGM approach consistently demonstrates superior or competitive performance compared to the alternative approaches. In terms of subgroup identification, SBJGM performs significantly better with much smaller CEs. Among the alternatives, the FM methods tend to exhibit the poorest subgroup identification performance, while SCAN and BCGS have better subgroup identification accuracy but are still much inferior compared to SBJGM.
@@
Figure 1: Simulation results for the power-law network under the scenarios with $K=2$ based on 100 replicates.
As the similarity between the networks increases (from S1 to S3), it is more challenging to identify subgroup memberships. However, SBJGM still shows more prominent advantages compared to SCAN, BCGS, FMGemBag, and FMJGL in terms of both heterogeneity and network identification performance. It is interesting that, under S3, although with a certain level of clustering errors, SBJGM has better network identification performance than TGemBag and TJGL with “ideal” true subgroup memberships, suggesting the effectiveness of the proposed similarity-based prior. Under the scenarios with an imbalanced design or a larger number of subgroups, performance of all approaches declines. Under these scenarios, the proposed approach, when compared to TGemBag, exhibits a slightly worse performance in network identification and estimation due to the more complex unknown heterogeneity and the relatively limited sample size. However, it still consistently outperforms the other alternatives. It maintains its superiority under the scenarios with various types of network structure, including the more challenging power-law networks with denser structures and hub structures. With relatively simpler nearest neighbor networks, when all methods generally perform well with higher TPR values, the proposed approach retains significantly fewer FPs compared to most alternatives.
@@
Figure 2: Simulation results for the Erdös-Rényi network under the scenarios with $K=2$ based on 100 replicates.
@@
Additionally, to gain a deeper understanding of the similarity-based prior, we illustrate a representative replicate under the scenario with setting S3 and an imbalanced design. The heatmaps depicting the true sparsity structures and the estimated results of the precision matrices with the proposed SBJGM, SCAN, and TGemBag, are presented in Figures S7-S8 (supplementary materials). The findings indicate that SBJGM excels in accurately identifying true sparsity structures and more effectively accommodating the similarity structures across different networks. To further evaluate uncertainty quantification, Figure S9 displays the PIPs versus the normalized true signals under the same setting. The estimated PIPs effectively capture uncertainty in network sparsity patterns, while a clear separation at the 0.5 threshold confirms the model’s accuracy in detecting true signals.
In addition, we conduct a series of extended simulation experiments, all examined under the power-law and S1 network structures. First, to evaluate performance under an increasing number of subgroups, we consider $K=4,5,6$, and $7$. Second, to investigate the impact of censoring, we examine scenarios with approximately $30\%,35\%$, $40\%$ and $45\%$ censoring rate under $K=2$. Third, to assess robustness to error distribution misspecification, we generate survival times from Weibull and log-logistic error distributions with $K=2$. The results, presented in Supplementary Figures S10-S12, consistently demonstrate that SBJGM accurately identifies subgroups and recovers network structures across these diverse scenarios.
@@
Following previous studies (Gao et al., 2016; Hao et al., 2018; Zhong et al., 2023), the main simulations assume that the number of subgroups $K$ is known for a more fair comparison. When $K$ is unknown, we recommend using the BIC for model selection. As an illustration, we consider scenarios with a power-law network and balanced design and evaluate the empirical performance of BIC over 100 replicates with candidate values $K=1,\ldots,5$. The selection frequencies, summarized in Table S2 (supplementary materials), indicate that BIC generally performs well. In particular, the S3 configuration, which exhibits greater subgroup similarity than S1 and S2, poses more challenges for accurate identification. For the BIC-selected values of $K$, Table S3 (supplementary material) reports the corresponding identification and estimation results. In cases where $K$ is incorrectly estimated, label alignment is performed minimizing the estimation error of $\boldsymbol{\Psi}$ (Ren et al., 2022). The results remain consistent with those presented in Figure 1 and Figure S4 (supplementary materials).